OpenStudy (anonymous):

In a small bag of M&Ms, there are 9 orange candies and 4 blue candies. You select 1 candy, eat it, and then select another. Find P(you select first an orange candy and then a blue candy). Round your answer to the nearest hundredth.

OpenStudy (anonymous):

8/33

OpenStudy (cwrw238):

P(ist orange) = 9/13 P(then select blue) = 4/12 reqd P = 9/13 * 4 / 12 = 9/13 * 1/3 = 9/39

OpenStudy (cwrw238):

or 0.23 to nearest hundredth

OpenStudy (anonymous):

thank you